10 <_2x<_x+9
x < 3x + 2 <2x + 6
x <_ 2x -1 <_x + 5
x<3x-1<2x+7
x - 10 <2(x-1)
(In the 2nd set of questions you solve the equations seperatley...according to my sheet)

Answer by venugopalramana(3286)

SEE THE FOLLOWING TO KNOW THE METHOD OF SOLUTION AND TRY..IF YOU STILL HAVE PROBLEM PLEASE COME BACK
2<5-2x/3<71 solutions --------------------------------------------------------------------------------Answer 8553 by venugopalramana(298) on 2005-10-30 11:25:19  2<5-2x/3<7..take one by one to solve2<5-2x/32x/3<5-2=3x<3*3/2x<9/25-2x/3<75-7<2x/3-2<2x/3...or..-2*3/2<1*x...or.....-3<1*x ....so the answer is -3<1*x<9/2********************************Graphs/17965: how do you solve eniqualities1 solutions --------------------------------------------------------------------------------Answer 8675 by venugopalramana(298) on 2005-11-01 11:13:46  the rules of transformations between equalities and inequalities are same but for few exceptions..they are given below..use them to solve the inequalities...........property..............equality....inequality.......conclusion...example........................3+2=5.......5<81.reverse sides.left to right ...5=3+2......8>5..inequality changes from < to > 2.multiply/divide with -ve number.-5=-(3+2)..-5>-8 same as above.........3.take reciprocal.................1/(3+2)=1/5..1/5>1/8..same as above..so take care of these and solve1example...x+5>8..or x>8-5=3..or x>32.3>5+x..or x+5<3..or x<3-5=-2..or x<-2..(exception rule no.1 above)2.example...-x+5>8..or-x>8-5=3..or -x>3..or x<3(exception rule no2. above)3.example...1/x>1/8...or x<8....(exception rule no.3 above)SO BE CAREFULL IN THESE OPERATIONS..PARTICULARLY MULTIPLICATION/DIVISION WITH UNKNOWNS..YOU HAVE TO BE SURE WHETHER IT IS A +VE NUMBER OR _VE NUMBER

Inequalities/18942: What is the solution to -4x <= -16 -4x is less than or equal to -16. how do i isolate the x from the -4x?1 solutions Answer 9138 by venugopalramana(345) on 2005-11-09 05:54:34  -4x <= -16let us cosider the 2 cases included here seperately1st.case is -4x=-16...............2nd.case is -4x<-16...........................1st.case is -4x=-16here on lhs we have - 4 multiplying x so when it is transfered to rhs it divides the number there.so we get x=-16/-4 =+42nd.case is -4x<-16here also on lhs we have - 4 multiplying x so when it is transfered to rhs it divides the number there.BUT IN CASE OF INEQUALITIES WHEN A -VE NUMBER IS USED TO MULTIPLY OR DIVIDE THEN THE INEQUALITY GETS REVERSED...FOR EXAMPLE WE KNOW5>0...BUT WHEN WE MULTIPLY BOTH SIDES WITH -1 SAY WE GET -5<0.THIS IS ONE OF THE FEW MAJOR DIFFERENCES WE HAVE BETWEEN INEQUALITIES AND EQUALITIESso we get x>-16/-4 ....or...x>+4hope you understood ..if you want to see more refer my solutions for similar problems posted in the site

Quadratic_Equations/18722: Could some one help me understand this problem? The instructions are to solve the inequality. State the solution set using interval notation and graph it:x - 2/x + 3 is less than 1 Thanks for the help1 solutions Answer 8977 by venugopalramana(345) on 2005-11-07 04:29:05  x - 2/x + 3 is less than 1 let y=(x - 2/x + 3)<1 x - 2/x + 3-1<0 ....or....x - 2/x + 2<0 (x^2-2+2x)/x<0...or ...(x^2+2x-2)/x<0..now a fraction will be -ve (<0)if n.r and d.r are of different signs..let us take the 2 cases dr=x is +ve...then nr should be -vedr=x is -ve...then nr should be +ve.now solve the nr using quadratic formula hence x=(-1+sqrt3)/2...and....(-1-sqrt3)/2..for conveinience if we call these 2 values as p(approximately=-1.37)and q(approximately=0.37),we find that if x lies between p and q (-1.37 and 0.37 )nr is -ve and when x is less than p(-1.37) or greater than q(0.37) ,nr is +ve.now we have to combine this with the above assumption on drdr=x is +ve...then nr should be -ve..so x should be between 0 and (-1+sqrt3)/2(not -1.37 to zero as x is already taken as positive)dr=x is -ve...then nr should be +ve.so x should be less than ....(-1-sqrt3)/2..(not 0.37 to zero as x is already taken as negative)